OpenStudy (dls):

A point moving with const. acc. from A to B, In a straight line AB has velocities u and v at A and B respectively.The velocity of point at C,the mid point of AB is?

OpenStudy (dls):

@mukushla :P

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

first find the distance AB using v^2 = u^2 + 2as s = (v^2 - u^2) / 2a = AB now we want the final velocity when displacement is s/2 substitute in the eq V^2 = u^2 + 2as and u'll get V = sqrt [ (u^2 + v^2)/2 ]

OpenStudy (anonymous):

@Vaidehi09 nice

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

ok m gonna take velocity at C as L. so, L^2 = u^2 + 2a ( s/2) L^2 = u^2 + 2a [ (v^2 - u^2) / 4a ] = u^2 + (v^2 - u^2)/2 = [ 2u^2 + v^2 - u^2 ] / 2 L^2 = [ u^2 + v^2 ] /2 so L = sqrt ( [u^2 + v^2]/2 )

OpenStudy (anonymous):

got it?

OpenStudy (dls):

got it thanks ! :)

OpenStudy (anonymous):

gr8!

OpenStudy (dls):

in this ques o.o if time from A to C is twice that from C to B then which of the following relation is correct? v=4u v=7u v=8u v=11u

OpenStudy (dls):

i tried v=u+at,ended up nowhere also the answer is v=7u :)

OpenStudy (anonymous):

i think i got it

OpenStudy (anonymous):

just give me a hint. not the whole sol.

OpenStudy (anonymous):

\[t_{AC}=\frac{2}{3} t_{AB}\]

OpenStudy (anonymous):

and \[v_{c}=\sqrt{\frac{u^2+v^2}{2}} \]

OpenStudy (anonymous):

i got v^2 +7u^2 -8uv = 0

OpenStudy (anonymous):

v^2 +7u^2 -8uv =(v-7u)(v-u)

OpenStudy (anonymous):

oh yea!...ok got it!

OpenStudy (anonymous):

thanks :))

OpenStudy (anonymous):

thank u too.